ok i no how to do algibra BUT!
i dont no when to add or subtract the varibles!
example 3x+6=8x-9 so..... do i add or subtract the 3x and 8x
please help me!!!!
from
hockey girl

Answer by checkley71(8403)

3X+6=8X-9 CONSIDER A VARIABLE A BANANA OR ANY OTHER OBJECT
8X-3X=6+9 YOU NOW HAVE 8(BANANAS) AND YOU SUBTRACT (TAKE AWAY) 3 BANANAS WHICH LEAVES YOU WITH 5 BANANAS OR 8X-3X=5X CONTINUING BY ADDING THE 6+9=15 WE GET
5X=15
NOW DIVIDE BOTH SIDESBY 5 AND YOU GET
5X/5=15/5 OR
X=3
PROOF
3*3+6=8*3-9
9+6=24-9
15=15